OpenStudy (anonymous):

Please Help! How do I solve this? (a^2^n - a^n - 6) / (a^n + 8).

OpenStudy (mathmath333):

\(\huge \tt \begin{align} \color{black}{ \dfrac{a^{2^{n} }- a^n - 6}{ a^n + 8} }\end{align}\) is it this

OpenStudy (anonymous):

a^n - 9 with remainder 66 It might help to make the simplifying substitution y = a^n, y^2 = a^2n, then do the division, then undo the substitution.

OpenStudy (anonymous):

Okay so can you show me in order the steps you did? So that I may learn how to complete other problems like this?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Think about rules for exponents. When you multiply powers of the same base, you add the exponents; likewise, when you divide powers, subtract the exponents, so ... a 2n a n =a 2n−n =a n .

OpenStudy (anonymous):

a^n will be the first term in the quotient, so multiply that by the divisor and subtract to find the new dividend. What is a^n × (a^n +8)?

OpenStudy (anonymous):

a^2n - a^n - 6a^n + 8a^n-(a^2n +8a^n)-9a^n - 6

OpenStudy (anonymous):

Ok so the answer is -16a^n-6?

OpenStudy (anonymous):

Or no?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok thank you very much.